I WOULD LIKE SOME HELP ON THE FOLLOWING QUESTION PLEASE.
Find the GCF.
36hk^3, 60k^2m, 84k^4n

Answer by Edwin McCravy(6936)


I WOULD LIKE SOME HELP ON THE FOLLOWING QUESTION PLEASE. 
Find the GCF. 
36hk^3, 60k^2m, 84k^4n

Break everything down into prime factors

36hk3 = 2·2·3·3·k·k·k

60k2m = 2·2·3·5·k·k·m

84k4n = 2·2·3·7·k·k·k·k·n

I notice that the first three factors 2·2·3 are common to 
all three expressions, so I will color them red.  You can 
just circle them on your paper:

36hk3 = 2·2·3·3·k·k·k

60k2m = 2·2·3·5·k·k·m

84k4n = 2·2·3·7·k·k·k·k·n

I can't color the remaining 3, the 5 or the 7 red because these 
are not common to all three expressions.  However, the first two 
k's are common to all three, so I can color them red: 

36hk3 = 2·2·3·3·k·k·k

60k2m = 2·2·3·5·k·k·m

84k4n = 2·2·3·7·k·k·k·k·n

I can't color any of the remaining k's in the first and third 
expressions red because they are not contained in the second.

So the GCF consists only of the red (or your circled) factors, so

GCF = 2·2·3·k·k 

Then multiply those together as

GCF = 12k2

If you do enough of these, you will learn to shorten this
process by doing some of it in your head.  You will also
notice that you will use the smallest exponent of any
letter that appears in all expressions.  But to get the
hang of GCF, do it this longer way for awhile.

Edwin